OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (3, 6) and (–2, 1)? y + 6 = 1(x + 3) y + 1 = –1(x + 2) y – 6 = –1(x + 3) y – 6 = 1(x – 3)

OpenStudy (anonymous):

d) y - 6 = 1(x-3)

OpenStudy (anonymous):

THANKS :)

OpenStudy (anonymous):

well to find the slope; it would be using the formula; y2 -y1 ------ x2 - x1 so plotting in those infos 1 - 6 ----- -2 - 3 -5 --- = 1 -5 so the slope is equal to 1; now the point slope form is y - y1 = m(x - x2) so it would be y - 6 = 1(x - 3) which would be d in your answer choices